# T HREAD : Thinking Deeper with Recursive Spawning

### v1

@@ Appendix E DataCommons QA and MIMIC-III ICU QA Benchmarks
We built the DataCommons QA benchmark utilizing data provided by Google Data Commons [9] with a focus on U.S. data. Data Commons contains publicly available data, aggregated from sources such as the Center for Disease Control and Prevention and the World Health Organization.
@@
Generation of the benchmark started with the production of questions which fit the following criteria: 1) answerable by information provided by Google Data Commons, 2) broad enough to encompass information sampled between different locations (eg. two counties), 3) lacking in subjectivity for the creation of a verifiable answer. The benchmark consists of question templates, which generate the questions of the benchmark, and ground-truth programs, which generate the answers to the questions of the benchmark. This format allows for the creation of a large quantity of diverse questions. Each question template consists of a sentence containing placeholders such as {variable_text} which, upon random sampling of a variable, are altered accordingly, facilitating the automated production of a variety of questions based on one question template. An example of a question template is as follows: ’Was {variable_text} in {entity1} increasing or decreasing from {time1} to {time2}?’. Time series data was sampled at the city, state, county, and country level for the United States. Data retrieval from Google Data Commons was conducted through the REST API provided by the website.
@@
A “ground-truth program” was created for each question template to generate a verifiably correct answer for each question within the benchmark. These programs varied in length from 100-200 lines of code and analyzed data acquired from the Google Data Commons API in accordance with the constraints tailored by the question it was answering. For example, if a question template included a constraint about the years of data to be sampled, the ground-truth program was constructed to generate a year range for the answer. Code within the program would subsequently be created to filter and analyze the data acquired from Google Data Commons accordingly by applying whichever mathematical concept was required for the answer (eg. correlation, linear regression, median, mean, etc.). Consequently, ground-truth programs formed the backbone of the benchmark through the creation of answers to the diverse questions which comprised the question template set. Figure 5 shows an example with the question “From 2015 to 2021, was the rate of asthma increasing faster in Boston or LA?”.
@@
The MIMIC-III ICU QA was created using the same approach as described above. However, questions were instead based on clinical time-series data provided by MIMIC-III [14]. We outline all steps, along with the code, to reproduce the benchmarks at https://github.com/philipmit/thread. We release the full dataset for DataCommons QA. Due to restrictions with MIMIC-III data access, we cannot directly release the dataset for MIMIC-III ICU QA. However, upon gaining access to MIMIC-III following the instructions outlined at https://physionet.org/content/mimiciii/1.4/, you can reproduce the full dataset using the code that we release.
@@
Appendix F Prompts
Below, we provide examples of the prompt used for each benchmark. The given task or question is highlighted in yellow. The text corresponding to the spawning of child threads is highlighted in blue, pink, orange, and green.
F.1 DataCommons QA
@@ F.2 MIMIC-III ICU QA
# The {variable} during {timeframe} for patient {patientID} was =>{data_patient_in_timeframe}<=
@@
ymean=np.mean(y)
print(ymean)
@@
table = pd.read_csv(PATIENTDATADIR, sep=‘’͡)
@@
filter_col_name = ‘MORTALITYINHOSPITAL’
@@
targetpatientgroup = table[table[variable_col_name]==1]
y = targetpatientgroup[variable_col_name]
ymean=np.mean(y)
print(ymean)
@@ F.4 WebShop
print(’You have retrieved {results_under_max_price}.’)
@@
print(’You have identified {item_to_purchase}.’)
@@
print(’This item has 1 attribute.’)
@@
print(’You have purchased {item_to_purchase}.’)
@@ F.5 TextCraft
print(’You have completed the action.’)
@@
print(’You cannot get the material.’)

### v2

@@ Appendix F DataCommons QA and MIMIC-III ICU QA Benchmarks
We built the DataCommons QA benchmark utilizing data provided by Google Data Commons Guha (2019) with a focus on U.S. data. Data Commons contains publicly available data, aggregated from sources such as the Center for Disease Control and Prevention and the World Health Organization.
@@
Generation of the benchmark started with the production of questions which fit the following criteria: 1) answerable by information provided by Google Data Commons, 2) broad enough to encompass information sampled between different locations (e.g., two counties), 3) lacking in subjectivity for the creation of a verifiable answer. The benchmark consists of question templates, which generate the questions of the benchmark, and ground-truth programs, which generate the answers to the questions of the benchmark. This format allows for the creation of a large quantity of diverse questions. Each question template consists of a sentence containing placeholders such as {variable_text} which, upon random sampling of a variable, are altered accordingly, facilitating the automated production of a variety of questions based on one question template. An example of a question template is as follows: ’Was {variable_text} in {entity1} increasing or decreasing from {time1} to {time2}?’. Time series data was sampled at the city, state, county, and country level for the United States. Data retrieval from Google Data Commons was conducted through the REST API provided by the website.
@@
A “ground-truth program” was created for each question template to generate a verifiably correct answer for each question within the benchmark. These programs varied in length from 100-200 lines of code and analyzed data acquired from the Google Data Commons API in accordance with the constraints tailored by the question it was answering. For example, if a question template included a constraint about the years of data to be sampled, the ground-truth program was constructed to generate a year range for the answer. Code within the program would subsequently be created to filter and analyze the data acquired from Google Data Commons accordingly by applying whichever mathematical concept was required for the answer (e.g., correlation, linear regression, median, mean). Consequently, ground-truth programs formed the backbone of the benchmark through the creation of answers to the diverse questions which comprised the question template set. Figure 8 shows an example with the question “From 2015 to 2021, was the rate of asthma increasing faster in Boston or LA?”.
@@
The MIMIC-III ICU QA was created using the same approach as described above. However, questions were instead based on clinical time-series data provided by MIMIC-III Johnson et al. (2016). We outline all steps, along with the code, to reproduce the benchmarks at https://github.com/philipmit/thread. We release the full dataset for DataCommons QA. Due to restrictions with MIMIC-III data access, we cannot directly release the dataset for MIMIC-III ICU QA. However, upon gaining access to MIMIC-III following the instructions outlined at https://physionet.org/content/mimiciii/1.4/, you can reproduce the full dataset using the code that we release.
@@
Appendix G Prompts
Below, we provide examples of the prompt used for each benchmark. The given task or question is highlighted in yellow. The text corresponding to the spawning of child threads is highlighted in blue, pink, orange, and green. The full prompts for each task can be seen at https://github.com/philipmit/thread.
G.1 DataCommons QA
@@ G.2 MIMIC-III ICU QA
# The {variable} during {timeframe} for patient {patient_ID} was =>{data_patient_in_timeframe}<=
@@
y_mean=np.mean(y)
print(y_mean)
@@
table = pd.read_csv(PATIENT_DATA_DIR, sep=‘’͡)
@@
filter_col_name = ‘MORTALITY_INHOSPITAL’
@@
target_patient_group = table[table[variable_col_name]==1]
y = target_patient_group[variable_col_name]
y_mean=np.mean(y)
print(y_mean)
@@ G.4 WebShop
print(‘You have retrieved {results_under_max_price}.’)
@@
print(‘You have identified {item_to_purchase}.’)
@@
print(‘This item has 1 attribute.’)
@@
print(‘You have purchased {item_to_purchase}.’)
@@ G.5 TextCraft
print(‘You have completed the action.’)
@@
print(‘You cannot get the material.’)